Karla's candle factory is considering two different machines. Machine A is highly automated with FC of $25,000 and VC of $.1/candle. Machine B is actually no machine but rather hand labor with FC of $10,000 and VC of $.5/candle. If demand for Karla's candles is 25,000, which machine should she pick?
Answer #1

If demand for Karla's candles is 25,000

Total Cost in Machine A = 25000 + 0.1*25000 = $ 27500

Total Cost in Machine B = 10000 + 0.5*25000 = $ 22500

Hence She Should pick Machine B for the above demand

Note:

Break Even Point where either machine can be purchase

Demand = (25000-10000)/(0.5-0.1) = 37500

Hence if Demand is equal to 37500 Unit any machine can be Purchase

  1. Demand = 37500, => Machijne A or Machine B can be purchased
  2. Demand >37500, => Machine A should be Purchased
  3. Demand < 37500 , => Machine B should be Purchased